A can is 6 inches tall with a 5 inch diameter. It is 75% full of water. What is the volume of the water ?
1
Expert's answer
2019-05-22T09:16:23-0400

Cylinder volume formula:


"V=R^2h\pi"

, where

"R^2\pi"

is base area.

Find the volume of water:


"V(water)=V*0.75=\pi*(5:2)^2*6*0.75=88.35"